A projectile with mass 0.50 kg is launched from an initial height of 2.4 meters and reaches a maximum height of 3.4 meters before landing on the ground (0.0 meters.) What is its change in gravitational potential energy from the beginning to the end?
Answer #1

iniital position = h1 = 2.4 m


initial potential energy = U1 = m*g*h1 = 0.5*9.8*2.4 = 11.76 J

final position = h2 = 0


final potential energy = U2 = m*g*h2 = 0


change in gravitational potential energy = dU = U1 - U2 = 11.76J